OpenStudy (anonymous):

**MEDAL/FAN A study of five hundred adults found that the number of hours they spend on social networking sites each week is normally distributed with a mean of 14 hours. The population standard deviation is 3 hours. What is the margin of error for a 95% confidence interval? 0.134 0.220 0.263 0.313

OpenStudy (anonymous):

@ganeshie8 @pooja195 @Michele_Laino

OpenStudy (michele_laino):

Hint: 95% confidence interval, means that our variable has to be within an interval whose ends are: \[\Large \begin{gathered} 14 - 1.96\sigma = 14 - 1.96 \times 3 = ...? \hfill \\ 14 + 1.96\sigma = 14 + 1.96 \times 3 = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

So its 8.12 and 19.88

OpenStudy (anonymous):

@Michele_Laino is that right? Can you help me pls thats not an option